# Safety of Multimodal Large Language Models on Images and Text

### v2

@@ Abstract
Attracted by the impressive power of Multimodal Large Language Models (MLLMs), the public is increasingly utilizing them to improve the efficiency of daily work. Nonetheless, the vulnerabilities of MLLMs to unsafe instructions bring huge safety risks when these models are deployed in real-world scenarios. In this paper, we systematically survey current efforts on the evaluation, attack, and defense of MLLMs’ safety on images and text. We begin with introducing the overview of MLLMs on images and text and understanding of safety, which helps researchers know the detailed scope of our survey. Then, we review the evaluation datasets and metrics for measuring the safety of MLLMs. Next, we comprehensively present attack and defense techniques related to MLLMs’ safety. Finally, we analyze several unsolved issues and discuss promising research directions. The latest papers are continually collected at https://github.com/isXinLiu/MLLM-Safety-Collection.
@@ 1 Introduction
We have witnessed the prosperous development of large language models (LLMs) in recent years, such as GPT-4222https://cdn.openai.com/papers/gpt-4.pdf, LLaMA-2333https://ai.meta.com/llama/ and Mixtral 8x7B [^4] https://mistral.ai/news/mixtral-of-experts/. The powerful capabilities of LLMs not only provide convenience for human life but also bring huge safety risks  Zhang et al. (2023); Sun et al. (2023). Much ink has been spent trying to make the LLMs safer by various alignment techniques (e.g.,  Rafailov et al. (2023)) and these methods have successfully enhanced the security of LLMs.
@@
Figure 2: Taxonomy: safety of MLLMs on images and text.
Based on such advancement of LLMs, researchers empower a LLM to handle other modalities beyond text via many multimodal fusion approaches. In this paper, we investigate Multimodal Large Language Models (MLLMs) on 2D images and text. If there is no additional explanation, the MLLM we refer to is assumed to be on 2D images and text.
MLLMs have drawn increasing attention due to their enormous multimodal potential (e.g., LLaVA [^5] https://github.com/haotian-liu/LLaVA, MiniGPT-4666https://github.com/Vision-CAIR/MiniGPT-4, Gemini [^7] https://blog.google/technology/ai/google-gemini-ai/, GPT-4V [^8] https://cdn.openai.com/papers/GPTV_System_Card.pdf). And there exist some works that design an attack to induce MLLMs to output unsafe content Fu et al. (2023); Bailey et al. (2023), propose a defense method Pi et al. (2024); Wu et al. (2023), or create a security evaluation dataset Chen et al. (2023b); Lin et al. (2024). However, compared to the research progress made in LLMs’ security, the study about MLLMs’ safety is still in its early stages. The lack of a comprehensive survey on MLLMs’ safety makes it hard to know the whole landscape of this field and we might wonder the following questions:
@@
  • How to measure (e.g., datasets, metrics) the safety level of MLLMs? A good evaluation technique is essential to tell how far MLLMs are from becoming safe.
@@
After conducting extensive research, we conclude that the risks from the image modality mainly include three aspects: (1) adding adversarial perturbations to images can bring satisfying attack results with low cost; (2) MLLMs based on aligned LLMs usually reject malicious textual instructions, but when leveraging the inherent Optical Character Recognition (OCR) ability they directly obey corresponding visual instructions; (3) cross-modal training weakens the alignment ability of aligned LLMs. To enable communities to understand MLLMs’ safety better, we present a survey to summarize the research progress from these perspectives: evaluation, attack, and defense (as shown in Figure 2). Our contributions are as follows:
@@
The rest of our survey unfolds as follows: Section 2 gives a brief overview of MLLMs and understanding of safety. Next, we sort out the benchmarks and metrics used for safety evaluation in Section 3. Then we describe attack techniques in Section 4 and introduce defense methods in Section 5. Finally, we discuss some future research directions in Section 6 and the paper concludes in Section 7.
@@ 2.1 An Overview of MLLMs
A MLLM [^9] This paper doesn’t consider tool-augmented LLMs and LLM agents. consists of a LLM ($\geq$1B), a vision encoder, and a cross-modal fusion module. After visual instructing tuning, the fusion module has learned cross-modal information, which helps LLMs handle image and text inputs and give proper responses in natural language. Current MLLMs can be divided into two categories: closed-source (e.g., GPT-4V, Bard [^10] https://bard.google.com/) and open-source MLLMs (e.g., LLaMA-Adapter V2111111https://github.com/OpenGVLab/LLaMA-Adapter, CogVLM [^12] https://github.com/THUDM/CogVLM, LLaVA-Phi Zhu et al. (2024)). For the latter, there are three common types of fusion modules: linear projection (e.g., LLaVA, MiniGPT-4, PandaGPT [^13] https://github.com/yxuansu/PandaGPT), learnable queries (e.g., InstructBLIP [^14] https://github.com/salesforce/LAVIS/tree/main/projects/instructblip, Qwen-vl [^15] https://github.com/QwenLM/Qwen-VL, BLIP-2161616https://github.com/salesforce/LAVIS/tree/main/projects/blip2), and cross-attention (e.g., IDEFICS [^17] https://huggingface.co/blog/idefics, OpenFlamingo [^18] https://github.com/mlfoundations/open_flamingo).
@@ 3.1 Dataset
Table 1: Comparison of recent representative evaluation datasets. The symbol $\ddagger$ denotes that “# Volume” only includes the multimodal test set that benchmarks the safety of MLLMs. The symbol *** denotes that ToViLaG Wang et al. (2023) only tests whether the image caption generated by a MLLM is toxic, which does not need ground truth captions.
Evaluation Dataset Data Source ${}^{\ddagger}$# Volume Evaluation Safety
@@
$*{}^{*}$oViLaG Wang et al. (2023) the NSFW dataset, - 21,559 Model-based Pornographic,
@@
[…]
@@ 3.2.2 Rule-based Evaluation
[…]
@@ 4 Attack
$*{}^{*}$ Bagdasaryan et al. (2023) Third party LLaVA, PandaGPT Targeted malicious text, Poisoned dialog
@@ 5 Defense
 Chen et al. (2023c) ${}^{\dagger}$${}_{ft}$ Reinforcement learning
@@ 6.3.1 Alignment Techniques
Reinforcement learning from human feedback (RLHF) is also a promising approach, which leverages human preferences as rewards to enable a model to align with human values and has played a huge role in the safety alignment of LLMs (e.g., InstructGPT [^26] https://github.com/openai/following-instructions-human-feedback). However, it is unclear what challenges will be encountered when applying RLHF to MLLMs for safety purposes. The construction of preference data is undoubtedly a crucial component and worth careful thinking from researchers.

### v3

@@ Abstract
Attracted by the impressive power of Multimodal Large Language Models (MLLMs), the public is increasingly utilizing them to improve the efficiency of daily work. Nonetheless, the vulnerabilities of MLLMs to unsafe instructions bring huge safety risks when these models are deployed in real-world scenarios. In this paper, we systematically survey current efforts on the evaluation, attack, and defense of MLLMs’ safety on images and text. We begin with introducing the overview of MLLMs on images and text and understanding of safety, which helps researchers know the detailed scope of our survey. Then, we review the evaluation datasets and metrics for measuring the safety of MLLMs. Next, we comprehensively present attack and defense techniques related to MLLMs’ safety. Finally, we analyze several unsolved issues and discuss promising research directions. The relevant papers are collected at https://github.com/isXinLiu/Awesome-MLLM-Safety.
@@ 1 Introduction
We have witnessed the prosperous development of large language models (LLMs) in recent years, such as GPT-4333https://cdn.openai.com/papers/gpt-4.pdf, LLaMA-2444https://ai.meta.com/llama/ and Mixtral 8x7B [^5] https://mistral.ai/news/mixtral-of-experts/. The powerful capabilities of LLMs not only provide convenience for human life but also bring huge safety risks  Zhang et al. (2023); Sun et al. (2023). Much ink has been spent trying to make the LLMs safer by various alignment techniques (e.g.,  Rafailov et al. (2023)) and these methods have successfully enhanced the security of LLMs.
@@
Figure 2: Taxonomy: safety of MLLMs on images and texts.
Based on such advancement of LLMs, researchers empower an LLM to handle other modalities beyond text via many multimodal fusion approaches. In this paper, we investigate Multimodal Large Language Models (MLLMs) on 2D images and text. If there is no additional explanation, the MLLM we refer to is assumed to be on 2D images and text.
MLLMs have drawn increasing attention due to their enormous multimodal potential (e.g., LLaVA [^6] https://github.com/haotian-liu/LLaVA, MiniGPT-4777https://github.com/Vision-CAIR/MiniGPT-4, Gemini [^8] https://blog.google/technology/ai/google-gemini-ai/, GPT-4V [^9] https://cdn.openai.com/papers/GPTV_System_Card.pdf). And there exist some works that design an attack to induce MLLMs to output unsafe content Fu et al. (2023); Bailey et al. (2023), propose a defense method Pi et al. (2024); Wu et al. (2023), or create a security evaluation dataset Chen et al. (2023b); Lin et al. (2024). However, compared to the research progress made in LLMs’ security, the study about MLLMs’ safety is still in its early stages. The lack of a comprehensive survey on MLLMs’ safety makes it hard to know the whole landscape of this field and we might wonder the following questions:
@@
  • How to measure (e.g., datasets, metrics) the safety level of MLLMs? A good evaluation technique is essential to tell how close MLLMs are to achieving safety.
@@
After conducting extensive research, we conclude that the risks from the visual modality mainly include three aspects: (1) adding adversarial perturbations to images can bring satisfying attack results with low cost; (2) MLLMs based on aligned LLMs usually reject malicious textual instructions, but when leveraging the inherent Optical Character Recognition (OCR) ability they directly obey corresponding visual instructions; (3) cross-modal training weakens the alignment ability of aligned LLMs. To enable communities to understand MLLMs’ safety better, we present a survey to summarize the research progress from these perspectives: evaluation, attack, and defense (as shown in Figure 2). Our contributions are as follows:
@@
The rest of our survey unfolds as follows: Section 2 gives a brief overview of MLLMs and understanding of safety. Next, we sort out the benchmarks and metrics used for safety evaluation in Section 3. Then we describe attack techniques in Section 4 and introduce defense methods in Section 5. Finally, we discuss some future research directions in Section 6, and the conclusion in Section 7.
@@ 2.1 An Overview of MLLMs
A MLLM [^10] This paper doesn’t consider tool-augmented LLMs and LLM agents. consists of a LLM ($\geq$1B), a vision encoder, and a cross-modal fusion module. After visual instructing tuning, the fusion module has learned cross-modal information, which helps LLMs handle image and text inputs and give proper responses in natural language. Current MLLMs can be divided into two categories: closed-source (e.g., GPT-4V, Bard [^11] https://bard.google.com/) and open-source MLLMs (e.g., LLaMA-Adapter V2121212https://github.com/OpenGVLab/LLaMA-Adapter, CogVLM [^13] https://github.com/THUDM/CogVLM, LLaVA-Phi Zhu et al. (2024)). For the latter, there are three common types of fusion modules: linear projection (e.g., LLaVA, MiniGPT-4, PandaGPT [^14] https://github.com/yxuansu/PandaGPT), learnable queries (e.g., InstructBLIP151515github.com/salesforce/LAVIS/tree/main/projects/instructblip, Qwen-vl [^16] https://github.com/QwenLM/Qwen-VL, BLIP-2171717https://github.com/salesforce/LAVIS/tree/main/projects/blip2), and cross-attention (e.g., IDEFICS [^18] https://huggingface.co/blog/idefics, OpenFlamingo [^19] https://github.com/mlfoundations/open_flamingo).
@@ 3.1 Dataset
Evaluation Dataset Data Source # Volume Evaluation Safety
@@
ToViLaG Wang et al. (2023) the NSFW dataset, - 21,559 Model-based Pornographic,
@@
Table 1: Comparison of recent representative evaluation datasets. The symbol $\ddagger$ denotes that “# Volume” only includes the multimodal test set that benchmarks the safety of MLLMs. The symbol * denotes that ToViLaG Wang et al. (2023) only tests whether the image caption generated by a MLLM is toxic, which does not need ground truth captions.
@@
(1) Visual prompts creation. SafeBench Gong et al. (2023) identifies 10 safety scenarios and queries GPT-4 to 50 unique malicious questions for each scenario. These questions are rephrased into imperative sentences and then transformed into visual prompts via typography. When giving a manually designed textual benign instruction and these visual prompts to MLLMs, SafeBench manually checks whether their responses follow the malicious visual prompts. Similar to SafeBench using typography, MM-SafetyBench Liu et al. (2023) transfers harmful key phrases from textual questions to images and prompts GPT-4Azure [^24] https://learn.microsoft.com/en-us/azure/ai-services/openai/ to measure MLLMs’ ability to discover the transfer.
@@ 3.2.2 Rule-based Evaluation
Some studies Bailey et al. (2023); Carlini et al. (2023); Luo et al. (2024) pre-define target strings and examine whether MLLMs’ outputs contain these strings. The target strings can be curse words Carlini et al. (2023), harmful instructions (e.g., “kidnap”Luo et al. (2024), malicious API calls Bailey et al. (2023) etc. In addition to these works,  Chen et al. (2023b); Lin et al. (2024); Van and Wu (2023) treat the evaluation process as classification tasks and apply relevant metrics (e.g., accuracy, F1). Concretely,  Chen et al. (2023b) explores MLLMs’ capability to balance utility and privacy protection. The former is measured by F1 and the latter is judged via protection score proposed by this work.  Lin et al. (2024); Van and Wu (2023) focus on meme-based social abuse detection, which is regarded as a binary classification task. While  Lin et al. (2024) reports the macro-averaged F1 and accuracy,  Van and Wu (2023) leverages AUROC and accuracy to assess whether MLLMs can detect abuse in multimodal memes. Not limited to discriminative tasks (e.g., classification),  Schlarmann and Hein (2023); Fu et al. (2023) also explore generative tasks (e.g., image captioning).  Schlarmann and Hein (2023) investigates malicious text and fake information generation on image captioning (CIDEr, BLEU-4) and visual question answering (accuracy) benchmarks.  Fu et al. (2023) uses Structural Similarity Index Measure (SSIM) to compare the similarity between an origin image and the perturbed adversarial image. This work also utilizes BLEU and Rouge to judge the utility of a response.
@@ 4 Attack
 Bagdasaryan et al. (2023) Third party LLaVA, PandaGPT Targeted malicious text, Poisoned dialog
@@ 5 Defense
 Chen et al. (2023c) DRESSft Reinforcement learning
@@
Table 3: Comparison of different defenses. Infer. Align.: inference-time alignment. Train. Align.: training-time alignment. The symbol $\dagger$ denotes that DRESSft is trained by  Chen et al. (2023c).
@@ 6.3.1 Alignment Techniques
Reinforcement learning from human feedback (RLHF) is also a promising approach, which leverages human preferences as rewards to enable a model to align with human values and has played a huge role in the safety alignment of LLMs. However, it is unclear what challenges will be encountered when applying RLHF to MLLMs for safety purposes. The construction of preference data is undoubtedly a crucial component and worth careful thinking from researchers.
@@ 7 Conclusion
Acknowledgements
This work was supported by the Joint Key Project (Project No. U23A20298) and Young Scientists Project (Project No. 62206097) of the National Natural Science Foundation of China, Shanghai Pujiang Talent Program (Project No. 22PJ1403000). One of the authors, Chao Yang, is supported by the Shanghai Post-doctoral Excellent Program (Grant No. 2022234). Any opinions expressed in this material are only those of the author(s).